OpenStudy (anonymous):

The distance of the points P(4,0) from the line x+ 2y -5 = 0 measured parallel to the line 2x+ 3y + 5 =0 is?

OpenStudy (anonymous):

@hartnn @ghazi @waterineyes

OpenStudy (anonymous):

Eq of the line parallel to line 2x+ 3y + 5 =0 passibg through P(4,0) is 2x +3 y - 8 =0

OpenStudy (anonymous):

Who is parallel??

OpenStudy (anonymous):

parallel to the line 2x+ 3y + 5 =0 is?

OpenStudy (tkhunny):

1) Parallel to 2x + 3y + 5 = 0 is a slope of -2/3. You need to see why this is. 2) Checking, just to see if it's easy. The slope of x + 2y - 5 = 0 is -1/2. Too bad. It's not perpendicular. 3) Point-Slope, then (y-0) = (-2/3)(x-4) = (-2/3)x + (8/3) 4) Find the point of intersection. 5) Measure the distance. There are other ways, depending on what material you are studying.

OpenStudy (anonymous):

I asked, two lines are parallel??

OpenStudy (tkhunny):

Two lines are not parallel. We are measuring the distance along a given line. We're just not quite given the line. That's what Step 3 is, above.